# Dependency-Aware Compilation for Surface Code Quantum Architectures

### v2

@@ Abstract.
Practical applications of quantum computing depend on fault-tolerant devices with error correction. Today, the most promising approach is a class of error-correcting codes called surface codes. We study the problem of compiling quantum circuits for quantum computers implementing surface codes. Optimal or near-optimal compilation is critical for both efficiency and correctness. The compilation problem requires (1) mapping circuit qubits to the device qubits and (2) routing execution paths between interacting qubits. We solve this problem efficiently and near-optimally with a novel algorithm that exploits the dependency structure of circuit operations to formulate discrete optimization problems that can be approximated via simulated annealing, a classic and simple algorithm. Our extensive evaluation shows that our approach is powerful and flexible for compiling realistic workloads.
@@ 1. Introduction
Quantum computation promises to surpass classical methods in important domains, potentially unlocking breakthroughs in materials science, chemistry, machine learning, and beyond. However, as individual physical qubits and operations are error-prone, these applications require an error-correction scheme for detecting and correcting faults. Quantum error-correction suppresses errors with redundancy: encoding the state of a single logical qubit using several physical qubits. Experimentalists have recently demonstrated error suppression for a single logical qubit (Zhao et al., 2022; Acharya et al., 2022; Ryan-Anderson et al., 2021) and small multi-qubit systems (da Silva et al., 2024; Bluvstein et al., 2023; Reichardt et al., 2024; Acharya et al., 2024).
To harness the full of the fault-tolerant quantum computers on the horizon, we need optimizing compilers that convert circuit-level descriptions of quantum programs to error-corrected elementary operations while preserving as much parallelism as possible. Quantum compute is a scarce resource, so inefficient compilation can be extremely costly. Further, the longer the computation, the higher the probability of logical errors, which affect the result.
@@
We target a well-studied type of error-correction scheme called a surface code (Knill et al., 1998; Fowler et al., 2012; Litinski, 2018). A surface code quantum device embeds logical qubits into a two-dimensional grid of physical qubits. Two-qubit gates impose limitations on the execution of a quantum circuit by introducing contention constraints. Each two-qubit gate occupies a path on the grid and simultaneous paths cannot cross. Gates which can theoretically be executed in parallel may be forced into sequential execution if the path of one “blocks” the other, as shown in Fig. 1. A compiler must carefully map qubits to grid locations and route two-qubit gates such that such conflicts between gates are minimized and parallelism is maximized. We call this the surface code mapping and routing (scmr) problem.
@@
Existing work on the scmr problem is limited along two axes: optimality and generality (see Table 1 for a summary): (1) optimality: some techniques do not optimize execution time (Watkins et al., 2023), or optimize routing with respect to a fixed, trivial mapping (Beverland et al., 2022a); (2) generality: other techniques (Hua et al., 2021; Javadi-Abhari et al., 2017; Zhu et al., 2023) do not account for routing t gates, a major source of two-qubit gates, and/or assume a sparse grid layout (Hua et al., 2021; Javadi-Abhari et al., 2017; Beverland et al., 2022a). This means they cannot be directly applied to other architectures which pack qubits into a smaller grid area (Litinski and Oppen, 2018; Watkins et al., 2023). Compact architectures are critical because they require fewer physical resources, but they make scmr more challenging.
In this paper, we present an optimizing mapping and routing algorithm that addresses the full generality of the scmr problem, targeting circuits with t gates and architectures with arbitrary layouts.
@@
Table 1. Feature summary of our approach and prior work
@@
Obtaining optimal or near-optimal scmr solutions is critical for both efficiency and correctness of quantum computation, but the problem is np-complete (see Appendix A for the proof). We find that brute-force solutions based on, e.g., satisfiability solvers, do not scale to realistic problem sizes, as previously shown in the analogous mapping and routing problem for near-term noisy devices without error correction (Wille et al., 2019; Tan and Cong, 2020; Lin et al., 2023; Molavi et al., 2022),
@@
Our goal with mapping is to map circuit qubits in such a way that is less likely to cause conflicts in routing gates. To do so, we need to track which pairs of qubits interact in the circuit. We extend this basic idea with the key insight that spatial relationships are only relevant between particular pairs of gates. For example, if one gate logically depends on another, then no valid routing solution will execute the two gates simultaneously, so there is no reason for the mapping stage to avoid crossing paths between them. By distinguishing between potential crossings which are likely to be problematic at the routing step and those which are not, we can better evaluate a candidate mapping. We capture this fine-grained information in a new data structure we call a layered interaction graph, which is the basis of our dependency-aware mapping approach.
@@
At the routing stage, our key insight is that the order in which gates are routed is crucial to the optimality of the scmr solution. Specifically, we formulate a discrete optimization problem over routing orders, with the goal of discovering a routing order that routes gates that are critical in the dependency structure of the circuit. A critical gate has lots of other gates that consume its results later in the circuit, so routing it early unlocks deeper layers of the circuit for routing. Highly critical gates are prioritized to avoid a long “tail” of unexploited parallelism as the gates on the critical path are executed in sequence. We observe that greedily routing gates in order of criticality, as considered in prior work (Javadi-Abhari et al., 2017), does not maximize the total criticality of routed gates because a single gate can block the execution of multiple gates of slightly lower criticality. However, by searching over the space of routing orders, we can find the order which is best with respect to total criticality.
@@
We also present a baseline algorithm that is guaranteed to provide an optimal solution to the scmr problem. The optimal algorithm is based on a novel sat encoding of scmr. Much like constraint-based approaches to mapping and routing for quantum computers without error correction (Tan and Cong, 2020; Lin et al., 2023; Shaik and van de Pol, 2024), we construct a boolean formula that is satisfiable if and only if there exists an scmr solution with a given execution time. Then, we find an optimal solution through a sequence of calls to a sat solver, starting from a known lower bound and incrementing until the formula is satisfiable. Though the sat-based approach does not scale to large-scale applications, it allows us to find exact solutions for small circuits and to precisely assess the optimality of dascot for these instances.
@@
  • A formalization of the surface code mapping and routing problem (scmr) capturing the combinatorial task of compiling circuits for surface code devices (Section 3).
  • A dependency-aware mapping algorithm using a new data structure called a layered interaction graph (Section 4.2)
  • A dependency-aware routing algorithm which iteratively chooses the next gates to route based on progress through the dependency structure of the circuit (Section 4.3)
  • A sat-based scmr algorithm for finding optimal solutions on small circuits (Section 5)
  • An empirical evaluation demonstrating the optimality and generality of dascot (Section 6)
@@ 2.2. An Overview of Our Approach
We design our approach to support arbitrary architectures. For example, given a 9 qubit circuit, Fig. 7 shows two standard architectures proposed in the literature (Litinski, 2018; Horsman et al., 2012; Watkins et al., 2023) which we call the Square Sparse and Compact architectures. These two architectures represent opposite ends of the space-time spectrum, with the Square Sparse architecture using many more qubits to allow routing more gates in parallel.
@@ 3. Surface Code Mapping & Routing
Fig. 8 shows a qubit map where four qubits are mapped to four of the 12 vertices in the architecture.
@@ 4. The dascot Algorithm
The scmr problem is np-complete (as shown in Appendix A), so exact optimization is infeasible for large instances. To simplify the problem, we decouple the mapping and routing and solve each separately. We present novel algorithms for mapping and routing that exploit circuit dependency structure and use simulated annealing to efficiently search the space of solutions. We start with a short primer on simulated annealing.
@@ 4.2. Dependency-Aware Mapping
In the mapping stage, our goal is to find a qubit map which is likely to admit a valid gate route with few steps without explicitly computing the full routing. We present an approach based on the idea of layered interaction graphs, which allows us to guide the search for a mapping using the dependency structure of a circuit.
@@ 4.3. Dependency-Aware Routing
At the routing stage, we take a circuit, architecture, and qubit map as input and return a sequence of time steps. As formalized in the definition of a valid gate route in Section 3, each time step $k$ consists of (1) a set of parallel gates, corresponding to the gates $g$ for which we set $R_{\textsc{time}}(g)=k$, and (2) disjoint paths along which to route them, defining the function $R_{\textsc{space}}:C\to\langle V\rangle$ on the same set of gates. The goal is to find the sequence with the fewest total time steps.
@@
We use a dependency-aware metric, where the cost reduction associated with routing a gate $g$ is equal to its criticality—we want to prioritize routing important gates first. The criticality of a gate is the depth of subcircuit of gates which depend on $g$, including $g$ itself. For example, in the circuit in Fig. 6, both gates have a criticality of 1, while the gate $\textsc{cnot}~{}q_{0}~{}q_{3}$ in Fig. 8 has a criticality of 3. Intuitively, routing critical gates unlocks deeper layers of the circuit for routing. Gates with high criticality should be prioritized to avoid a long “tail” of underutilized time steps as the gates on the critical path are executed in sequence.
@@ 5. A sat-Based Optimal Baseline
5.1. Constraints
@@ 6. Implementation and Evaluation
Our evaluation is performed on a benchmark suite of 232 application circuits. Our suite includes the entire set collected by Zulehner et al. (2018). This existing set consists of circuits derived from the RevLib suite (Wille et al., 2008) and programs written in the Quipper (Green et al., 2013) and ScaffoldCC (JavadiAbhari et al., 2014) quantum programming languages. We extended it with implementations of major quantum algorithms: Shor’s Algorithm (Shor, 1994), the Quantum Fourier Transform (Coppersmith, 2002), Bernstein-Vazirani (Bernstein and Vazirani, 1997), QAOA (Farhi et al., 2014), and Grover’s Algorithm (Grover, 1996).
@@
We target the two architectures Square Sparse and Compact from Fig. 7. The Square Sparse architecture represents a case with a large space footprint studied in prior work on mapping and routing (Hua et al., 2021; Watkins et al., 2023). It is the smallest square grid which includes enough space to surround each circuit qubit with routing ancillae on all sides. For a circuit with $n$ qubits, this results in a side length of $12\lceil\sqrt{n}\rceil+$ 1. The Compact architecture, in contrast, is designed to minimize the required device qubits. This architecture is nearly linear, with three rows, and enough columns to include an ancilla qubit between each circuit qubit ($n-1$ columns for even $n$). In both cases, we assume this mapping region is surrounded on all sides by magic state qubits.
@@
(RQ1) Comparison to prior work
@@
On the Square Sparse architecture, the largest category by far is “Match” where the two algorithms reach equivalent solutions, containing 83% of the circuits. This is generally because, with the plentiful routing resources of this architecture, both algorithms achieve the theoretical lower-bound given by the depth of the circuit. On the other hand, the constraints of the Compact architecture do reveal a separation between the two algorithms. dascot outperforms or matches Autobraid on 84% of the benchmarks and strictly outperforms Autobraid on 59% of the benchmarks.
@@
First, to assess our mapping algorithm, we compare against a version of our algorithm called dascot-randmap which chooses a qubit map at random, then applies the dascot routing algorithm (Alg. 3). Choosing a random map is equivalent to setting the initial temperature of the simulated annealing search equal to the final temperature, performing zero iterations of the search, or not applying a mapping optimization pass at all as in the lattice surgery compiler (LSC) (Watkins et al., 2023).
@@
We see more uniform cost reduction for the routing phase. We analogously define the dascot-randroute algorithm which first applies the dascot mapping algorithm (Alg. 1). Then, for routing, it replaces the minimization procedure with a random order selection, once again corresponding to zero iterations of simulated annealing and the procedure applied by LSC (Watkins et al., 2023). In Fig. 17(middle), we compare dascot to dascot-randroute. The results indicate that a dependency-aware analysis of routing order leads to significantly stronger solutions than committing to a single random order at each step. We see a strict improvement in mean cost ratio on 90% of application circuits from applying dependency-aware routing for a mean relative improvement of 6%. Similar findings hold when we compare dascot to choosing a routing order by sorting gates according to routing path length and criticality. We show the results for the shortest-first routing order, which is the approach of EDPC (Beverland et al., 2022a), in Fig. 17(right). The plot for the critical-first sorting order closely resembles this one and is included in Appendix D.
@@
(RQ4) Optimality-Scalability tradeoff

### v3

@@ Abstract.
Practical applications of quantum computing depend on fault-tolerant devices with error correction. Today, the most promising approach is a class of error-correcting codes called surface codes. We study the problem of compiling quantum circuits for quantum computers implementing surface codes. Optimal or near-optimal compilation is critical for both efficiency and correctness.
The compilation problem requires (1) mapping circuit qubits to the device qubits and (2) routing execution paths between interacting qubits. We solve this problem efficiently and near-optimally with a novel algorithm that exploits the dependency structure of circuit operations to formulate discrete optimization problems that can be approximated via simulated annealing, a classic and simple algorithm. Our extensive evaluation shows that our approach is powerful and flexible for compiling realistic workloads.
quantum error-correction, simulated annealing
copyright: ccdoi: 10.1145/3720416journalyear: 2025journal: PACMPLjournalvolume: 9journalnumber: OOPSLA1article: 82publicationmonth: 4ccs: Software and its engineering Compilersccs: Hardware Quantum computation
@@ 1. Introduction
Quantum computation promises to surpass classical methods in important domains, potentially unlocking breakthroughs in materials science, chemistry, machine learning, and beyond. However, as individual physical qubits and operations are error-prone, these applications require an error-correction scheme for detecting and correcting faults. Quantum error-correction suppresses errors with redundancy: encoding the state of a single logical qubit using several physical qubits. Experimentalists have recently demonstrated error suppression for a single logical qubit (Zhao et al., 2022; Google Quantum AI, 2023; Ryan-Anderson et al., 2021) and small multi-qubit systems (da Silva et al., 2024; Bluvstein et al., 2023; Reichardt et al., 2024; Google Quantum AI and Collaborators, 2024).
To harness the full power of the fault-tolerant quantum computers on the horizon, we need optimizing compilers that convert circuit-level descriptions of quantum programs to error-corrected elementary operations while preserving as much parallelism as possible. Quantum compute is a scarce resource, so inefficient compilation can be extremely costly. Further, the longer the computation, the higher the probability of logical errors, which affect the result.
@@
We target a well-studied type of error-correction scheme called a surface code (Knill et al., 1998; Fowler et al., 2012; Litinski, 2019). A surface code quantum device embeds logical qubits into a two-dimensional grid of physical qubits. Two-qubit gates impose limitations on the execution of a quantum circuit by introducing contention constraints. Each two-qubit gate occupies a path on the grid and simultaneous paths cannot cross. Gates which can theoretically be executed in parallel may be forced into sequential execution if the path of one “blocks” the other, as shown in Fig. 1. A compiler must carefully map qubits to grid locations and route two-qubit gates such that conflicts between gates are minimized and parallelism is maximized. We call this the surface code mapping and routing (scmr) problem.
@@
Existing work on the scmr problem is limited along two axes, optimality and generality (see Table 1 for a summary): (1) optimality: some techniques do not optimize execution time (Watkins et al., 2024), or optimize routing with respect to a fixed, trivial mapping (Beverland et al., 2022a); (2) generality: other techniques (Hua et al., 2021; Javadi-Abhari et al., 2017; Zhu et al., 2024) do not account for routing t gates, a major source of two-qubit gates, and/or assume a sparse grid layout (Hua et al., 2021; Javadi-Abhari et al., 2017; Beverland et al., 2022a). This means they cannot be directly applied to other architectures which pack qubits into a smaller grid area (Litinski and Oppen, 2018; Watkins et al., 2024). Compact architectures are critical because they require fewer physical resources, but they make scmr more challenging.
We present an optimizing mapping and routing algorithm that addresses the full generality of the scmr problem, targeting circuits with t gates and architectures with arbitrary layouts.
Table 1. Feature summary of our approach and prior work
@@
Obtaining optimal or near-optimal scmr solutions is critical for both efficiency and correctness of quantum computation, but the problem is np-complete (Zhu et al., 2024). We find that brute-force solutions based on, e.g., satisfiability solvers, do not scale to realistic problem sizes, as previously shown in the analogous mapping and routing problem for near-term noisy devices without error correction (Wille et al., 2019; Tan and Cong, 2020; Lin et al., 2023; Molavi et al., 2022),
@@
Our goal with mapping is to map circuit qubits in such a way that is less likely to cause conflicts in routing gates. To do so, we need to track which pairs of qubits interact in the circuit. While this idea has appeared in different guises in the qubit mapping literature, existing solutions tend to be flow-insensitive in nature (Javadi-Abhari et al., 2017; Hua et al., 2021), meaning that qubit interactions are modeled at a rough granularity that does not take into account gate ordering in the circuit.
We make the key insight that to build a good candidate mapping, we need finer-grained, flow-sensitive modeling of qubit interactions. For example, if one gate logically depends on another, then no valid routing solution will execute the two gates simultaneously, so there is no reason for the mapping stage to avoid crossing paths between them. By distinguishing between potential crossings which are likely to be problematic at the routing step and those which are not, we can better evaluate a candidate mapping. We capture this information in a new data structure we call a layered interaction graph, which is the basis of our dependency-aware mapping approach.
@@
At the routing stage, our key insight is that the order in which gates are routed is crucial to the optimality of the scmr solution. Specifically, we formulate a discrete optimization problem over routing orders, with the goal of discovering a routing order that routes gates that are critical in the dependency structure of the circuit. A critical gate has lots of other gates that consume its results later in the circuit, so routing it early unlocks deeper layers of the circuit for routing. Highly critical gates are prioritized to avoid a long “tail” of unexploited parallelism as the gates on the critical path are executed in sequence. We observe that greedily routing gates in order of criticality, as considered in prior work (Javadi-Abhari et al., 2017), does not maximize the total criticality of routed gates because a single gate can block the execution of multiple gates of slightly lower criticality. However, our new perspective on the routing problem as a search over the space of routing orders allows us to find the order which is best with respect to total criticality.
@@
We present an optimal algorithm for the scmr problem based on a novel sat encoding of scmr. Much like constraint-based approaches to mapping and routing for quantum computers without error correction (Tan and Cong, 2020; Lin et al., 2023; Shaik and van de Pol, 2024), we construct a boolean formula that is satisfiable if and only if there exists an scmr solution with a given execution time. Then, we find an optimal solution through a sequence of calls to a sat solver, starting from a known lower bound and incrementing until the formula is satisfiable. Though the sat-based approach does not scale to large-scale applications, it allows us to find exact solutions for small circuits and to precisely assess the optimality of dascot for these instances.
@@
  • A formalization of the surface code mapping and routing problem (scmr) capturing the combinatorial task of compiling circuits for surface code devices (Section 3). Compared to existing work, our formalization models the problem in more detail and with more flexibility, explicitly modeling magic states and allowing for arbitrary grids.
  • A dependency-aware mapping algorithm using a new data structure called a layered interaction graph (Section 4.2). In comparison to prior work (Javadi-Abhari et al., 2017; Hua et al., 2021), the layered nature of our interaction graphs captures finer-grained—flow-sensitive—interaction patterns between qubits, allowing us to choose better candidate mappings that maximize parallelism.
  • We make the key observation is that routing order is crucial for the optimality of scmr solutions. We therefore present a dependency-aware routing algorithm which searches over routing orders and chooses the next gates to route based on progress through the dependency structure of the circuit (Section 4.3). This is in contrast to greedy approaches (Hua et al., 2021) that choose to route gates by order of their importance, disregarding the downstream effects of such actions, and end up with suboptimal global solutions.
  • A sat-based scmr algorithm for finding optimal solutions on small circuits (Section 5). To our knowledge, this is the first optimal solver for the mapping and routing problem on a fault-tolerant architecture. The sat-based algorithm allows us to quantify the distance from optimality for our approximate, simulated-annealing-based solutions.
  • An empirical evaluation demonstrating the optimality and generality of dascot (Section 6).
@@ 2.2. An Overview of Our Approach
We design our approach to support arbitrary architectures. For example, given a 9 qubit circuit, Fig. 7 shows two standard architectures proposed in the literature (Litinski, 2019; Horsman et al., 2012; Watkins et al., 2024) which we call the Square Sparse and Compact architectures. These two architectures represent opposite ends of the space-time spectrum, with the Square Sparse architecture using many more qubits to allow routing more gates in parallel.
@@ 3. Surface Code Mapping & Routing
Fig. 8 shows a qubit map where four qubits are mapped to four of the 15 vertices in the architecture.
@@ 4. The dascot Algorithm
The scmr problem is np-complete (Zhu et al., 2024), so exact optimization is infeasible for large instances. To simplify the problem, we decouple the mapping and routing and solve each separately. We present novel algorithms for mapping and routing that exploit circuit dependency structure and use simulated annealing to efficiently search the space of solutions. We start with a short primer on simulated annealing.
@@ 4.2. Dependency-Aware Mapping
In the mapping stage, our goal is to find a qubit map which is likely to admit a valid gate route with few steps without explicitly computing the full routing. We present an approach based on a novel data structure called a layered interaction graph. Whereas prior work takes a “flattened” view of qubit interactions with no consideration of the ordering of gates, layered interaction graphs guide our search for a mapping using the dependency structure of a circuit.
@@ 4.3. Dependency-Aware Routing
At the routing stage, we take a circuit, architecture, and qubit map as input and return a sequence of time steps. As formalized in the definition of a valid gate route in Section 3, each time step $k$ consists of (1) a set of parallel gates, corresponding to the gates $g$ for which we set $R_{\textsc{time}}(g)=k$, and (2) disjoint paths along which to route them, defining the function $R_{\textsc{space}}:C\to\langle V\rangle$ on the same set of gates. The goal is to find the sequence with the fewest total time steps. The distinguishing feature of our approach is that we define and solve a discrete optimization problem over the space of candidates for the each time step, focusing on gates with high criticality.
@@
We use a dependency-aware metric, where the cost reduction associated with routing a gate $g$ is equal to its criticality—we want to prioritize routing important gates first.
Criticality: Let $C$ be a quantum circuit and $g$ be a gate in $C$. The criticality of $g$, denoted $\operatorname*{\textsc{criticality}}(g)$ is the depth of the subcircuit of $C$ consisting of gates which depend on $g$, including $g$ itself.
For example, in the circuit in Fig. 6, both gates have a criticality of 1, while the gate $\textsc{cnot}~{}q_{0}~{}q_{3}$ in Fig. 8 has a criticality of 3. Intuitively, routing critical gates unlocks deeper layers of the circuit for routing. Gates with high criticality should be prioritized to avoid a long “tail” of underutilized time steps as the gates on the critical path are executed in sequence.
@@ 5. A sat-Based Optimal Baseline
Next, we define the constraints over these variables that ensure a satisfying assignment corresponds to a mapping and routing solution.
@@ 6. Implementation and Evaluation
Our evaluation is performed on a benchmark suite of 232 application circuits. Our suite includes the entire set collected by Zulehner et al. (2018). This existing set consists of circuits derived from the RevLib suite (Wille et al., 2008) and programs written in the Quipper (Green et al., 2013) and ScaffoldCC (Javadi-Abhari et al., 2014) quantum programming languages. We extended it with implementations of major quantum algorithms: Shor’s Algorithm (Shor, 1994), the Quantum Fourier Transform (Coppersmith, 2002), Bernstein-Vazirani (Bernstein and Vazirani, 1997), QAOA (Farhi et al., 2014), and Grover’s Algorithm (Grover, 1996).
@@
We target the two architectures Square Sparse and Compact from Fig. 7. The Square Sparse architecture represents a case with a large space footprint studied in prior work on mapping and routing (Hua et al., 2021; Watkins et al., 2024). It is the smallest square grid which includes enough space to surround each circuit qubit with routing ancillae on all sides. For a circuit with $n$ qubits, this results in a side length of $12\lceil\sqrt{n}\rceil+$ 1. The Compact architecture, in contrast, is designed to minimize the required device qubits. This architecture is nearly linear, with three rows, and enough columns to include an ancilla qubit between each circuit qubit ($n-1$ columns for even $n$). In both cases, we assume this mapping region is surrounded on all sides by magic state qubits.
@@
6.1. (RQ1) Comparison to Prior Work
@@
On the Square Sparse architecture, the largest category by far is “Match” where the two algorithms reach equivalent solutions, containing 83% of the benchmark suite. In comparison, the “dascot better” category contains 16% of circuits and “Autobraid better” category contains 1% (just three circuits). Most circuits are in the “Match” category because, with the plentiful routing resources of this architecture, both algorithms often achieve the theoretical lower-bound given by the depth of the circuit. On the other hand, the constraints of the Compact architecture do reveal a separation between the two algorithms. dascot outperforms or matches Autobraid on 84% of the benchmarks and strictly outperforms Autobraid on 59% of the benchmarks.
@@
First, to assess our mapping algorithm, we compare against a version of our algorithm called dascot-randmap which chooses a qubit map at random, then applies the dascot routing algorithm (Alg. 3). Choosing a random map is equivalent to setting the initial temperature of the simulated annealing search equal to the final temperature, performing zero iterations of the search, or not applying a mapping optimization pass at all as in the lattice surgery compiler (LSC) (Watkins et al., 2024).
@@
We see more uniform cost reduction for the routing phase. We analogously define the dascot-randroute algorithm which first applies the dascot mapping algorithm (Alg. 1). Then, for routing, it replaces the minimization procedure with a random order selection, once again corresponding to zero iterations of simulated annealing and the procedure applied by LSC (Watkins et al., 2024). In Fig. 17(middle), we compare dascot to dascot-randroute. The results indicate that a dependency-aware analysis of routing order leads to significantly stronger solutions than committing to a single random order at each step. We see a strict improvement in mean cost ratio on 90% of application circuits from applying dependency-aware routing for a mean relative improvement of 6%. Similar findings hold when we compare dascot to choosing a routing order by sorting gates according to routing path length and criticality. We show the results for the shortest-first routing order, which is the approach of EDPC (Beverland et al., 2022a), in Fig. 17(right). The plot for the critical-first sorting order closely resembles this one and is included in Appendix D.
@@
6.4. (RQ4) Optimality-Scalability Tradeoff